OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 2x+y=5 Y+3=2x Here are the choices: 2x+2x-3=5 2x+3-2x=5 2x+2x+3=5

OpenStudy (radar):

O.K. The 2nd equation is 2x = y + 3 we can substitute for the 2x in the 1st equation: (y + 3) + y = 5 Now it is just algebra 2y + 3 = 5 after combining similar terms 2y = 5 - 3 subtracting three from both side. Please continue with the solution.